# Generative ensemble deep learning severe weather prediction from a deterministic convection-allowing model

### v1

@@ 1 Introduction
Validating the probabilistic severe weather predictions and exploring the advances of deep generative models are the main motivations of this work. Ensemble predictions have clear benefits. The best guess of an ensemble is typically more skillful than deterministic predictions. The ensemble spread measures the predictive uncertainties; it answers the question of “how much can we trust the predicted results?”. Given the difficulty and social impact of severe weather post-processing, organizing it within an ensemble prediction workflow is essential. Many deterministic, point-based machine learning methods can make ensemble predictions if they have an ensemble of CAM runs as inputs. However, maintaining and pre-processing a large ensemble of CAMs may introduce technical challenges such as data transmission and storage. The development of deep generative models provides an alternative. In this research, deep generative models are trained to create an ensemble of synthetic forecasts from a deterministic CAM run. This synthetic ensemble will provide convective-scale information and support the estimation of severe weather. Deep generative models are neural networks that can approximate probability distributions, and notably, they can create synthetic outputs from the data distributions they have learned (Creswell et al., 2018; Goodfellow et al., 2014; Ruthotto and Haber, 2021). Deep generative models have achieved success in downscaling [e.g., Leinonen et al. (2020) for satellite imagery, Miralles et al. (2022) for surface wind] and nowcasting problems [e.g., Ravuri et al. (2021), Luo et al. (2022), and Tian et al. (2019) for precipitation nowcasting, Gong et al. (2023) for radar reflectivity], indicating that they can be applied to generate high-resolution meteorological fields.
In this research, a novel ML-based CAM post-processing method is introduced that combines a CNN-based prediction model and a deep generative model that creates synthetic CAM forecasts as severe weather prediction inputs. The generative model can effectively generate an ensemble of severe weather predictions from a deterministic CAM run, potentially improving forecast skills and uncertainty estimates. The post-processing system uses CAM output on the native model grid, eliminating the need for upscaling that has been used in prior work (e.g., Loken et al. (2020); Sobash et al. (2020)). The system is trained to output probabilities of severe weather at grid points within the CAM domain using observations of severe weather events.
@@ 2.2 Forecast data
We use both HRRR version 3 and 4 (Dowell et al., 2022). The HRRR is an operational, real-time CAM based on WRF-ARW [Weather Research and Forecasting model, WRF; Advanced Research WRF, ARW; Skamarock et al. (2019)]. HRRR has 3 km horizontal grid spacing and 50 vertical levels; it takes Rapid Refresh (RAP; Benjamin et al. (2016)) as initial and boundary conditions, producing hourly, deterministic forecasts over CONUS. Here, the $0$ UTC initializations are used, with HRRRv3 covering 12 July 2018 to 2 December 2020, and HRRRv4 afterwards. In addition, experimental HRRRv4 initialized from 1 October 2019 to 2 December 2020 was also used.
@@ 2.3 Observations
Severe weather reports collected by the SPC at the National Oceanic and Atmospheric Administration (hereafter, “SPC reports”) are used as the observational target. SPC reports provide the location and time for tornadoes, hail, and convective wind gusts that have sufficient intensities and societal impact. Such information is placed on the gridded 80-km CONUS domain by matching the closest grid cell to the starting location of each report. For temporal dimensions, SPC reports are aggregated hourly with a 4-hour time window. That said, given a severe weather report on hour $t$; it will be mapped to the 80-km CONUS domain four times, on $t-2$, $t-1$, $t$, and $t+1$. This choice is based on the operational demand for 4-hour severe weather outlooks (e.g. Krocak and Brooks, 2020).
Gridded SPC reports are used for the training and verification of the post-processing models, as well as the estimation of climatology references (Fig. 1.b and d). We prefer the gridded SPC reports over the original ones because the spatiotemporal re-gridding creates a “tolerance” for the regional and population bias that SPC reports may potentially have (e.g. Doswell et al., 1999). Hereafter, when “SPC reports” is mentioned, it means the re-gridded version on the 80-km CONUS domain.
@@ 3 Methods
The above post-processing method was trained from 5 July 2018 to 31 December 2020 with HRRRv3 and HRRRv4 forecasts as inputs and SPC reports as targets. It produces 4-hr, 80-km severe weather probabilities out to 24 hours with the exceptions of 00Z and 01Z; these two forecast lead times are ignored because of the HRRR model spin-up. The verification period of the final post-processing outputs is 1 January 2021–31 December 2021 and with HRRRv4 only. The hyperparameters of the CGANs and the CNN-based prediction systems are adjusted on a trial-and-error basis. Neural networks with different combinations of learning rates, sample rebalancing strategies, and network sizes are trained and compared on the same validation set. The neural network with the best-performing combination is selected and applied in the inference stage.
@@ 3.1 Conditional Generative Adversarial Networks
Figure 3: (a) The training procedure of CGAN. (b) The architecture of the CNN-based generator; it contains convolutional layers (“conv”), Gaussian Error Linear Unit (GELU) activation function, and Batch Normalization (BN). Numbers beside each block represent the number of convolution kernels. (c) similar to (b), but for the architecture of the discriminator. Note that the CNN-based generator (b) will be preserved after training, whereas the discriminator is used for the CGAN training only.
This research applies CGANs to generate synthetic ensembles from deterministic HRRR forecasts. CGANs are Generative Adversarial Networks (GANs) that utilize conditional information as additional inputs to enable more control over the sample generation process (Mirza and Osindero, 2014; Isola et al., 2017). Here the conditional information is provided by five storm-scale explicit predictors: 2-5 km UH, 0-2 km UH, APCP, GRPL, and 10-m SPD. By accepting such conditional inputs, CGANs are trained to generate synthetic CREF and environmental predictors. The idea behind this choice is that UH and CREF are the two major CAM outputs that are related to convective storms directly (Clark et al., 2012; Gallo et al., 2021). If both UH and CREF are allowed to be generated by CGANs freely, the resulting synthetic ensembles may contain unrealistic storms and bias the severe weather prediction step. Thus, training CGANs to create CREF, but letting the generation process be conditioned on UH and other explicit predictors would balance the creativity and realism of CGAN outputs.
The CGAN in this research consists of a CNN-based generator and discriminator (Fig. 3a). The generator has two input branches, one takes conditional inputs, and the other accepts the initial state of predictors to be generated. Such initial states are prepared by combining the original HRRR predictors with $\mathcal{N}(0,0.5)$ distributed Gaussian noise (Fig. 3b). Negative values will be corrected for the initial states of CAPE, CIN, and CREF. The base architecture of the generator is an encoder-decoder CNN with skip connections on each encoding level, known as UNET (Ronneberger et al., 2015). The UNET applies convolutional layers with two strides and transposed convolutional layers for down- and up-samplings, respectively. The number of convolution kernels doubles after each down-sampling operation, from 32 to 512, and the numbers are symmetrical on the upsampling side (Fig. 3b). The discriminator of the CGAN is a CNN-based binary classifier that aims to distinguish generator outputs from their corresponding HRRR predictors. Besides the classification inputs, the discriminator also takes the same conditional inputs as that of the generator (Fig. 3c). For both CNNs, Batch Normalization (BN; Ioffe and Szegedy (2015)) is applied after each convolutional layer, and Gaussian Error Linear Unit (GELU; Hendrycks and Gimpel (2016)) is the activation function.
@@
Two CGANs of the above are proposed: one generates CAPE, CIN, and CREF (i.e., three input channels), and the other generates MSLP, 2-m Temp, 2-m Dewpoint T, 0-1 km SRH, 0-3 km SRH, 0-6 km U Shear, 0-6 km V Shear (i.e., seven input channels). The same CGANs are applied to all forecast lead times. The two CGANs are trained for 200 epochs using the Adaptive moment estimation (Adam) optimizer (Kingma and Ba, 2014) and learning rates of 1E-4. After training, CGAN generators are preserved and carried to the next post-processing stage.
@@ 3.2 CNN-based end-to-end severe weather prediction
Figure 4: (a) The CNN-based end-to-end severe weather prediction model. (b) the architecture of the 2-d CNN in (a); it contains convolutional layers (“conv”), Gaussian Error Linear Unit (GELU) activations, and Batch Normalization (BN). (c) The architecture of the classification model in (a). (d) The architecture of the MLP-based severe weather prediction model. Monte Carlo (MC) dropout layers are included in (c) and (d)
@@
The representation learning model is trained from 5 July 2018 to 31 December 2020, using HRRRv3 and v4 forecasts. SPC reports on the corresponding forecast lead times were used as training target. This is achieved by configuring an additional dense layer with a sigmoid activation function to compute cross-entropy loss (Fig. 4b). The training was based on the sample rebalancing of 1:10 for severe and non-severe cases. The classification model was trained using the same samples as the representation learning model, but independently for each forecast lead time window and with 1:1 sample rebalancing. The deactivation rate of MC dropout layers is 0.1. Both models use the Adam optimizer with initial learning rates of 1E-4, learning rate decay, and early stopping.
3.3 Post-processing experiments and baseline methods
@@
The MLP ensemble does not rely on CNNs to make severe weather predictions; instead, it combines an MLP with feature engineering steps. Given the same 64-by-64-sized HRRR predictors that CNN-based methods would use, the MLP baseline converts them into scalar inputs by computing their mean and maximum values on spatial and forecast lead time dimensions. This pre-processing step is similar to Sobash et al. (2020). The architecture of MLP is illustrated in Fig. 4d; it consists of two hidden layers with BN, Rectified Linear Unit (ReLU) activation function, and MC dropout. Several existing works have shown that MLPs are effective in deriving severe weather probabilities from CAM outputs (e.g. Sobash et al., 2020). Notably, they can outperform the conventionally used surrogate methods. Thus, in this experiment, the MLP baseline is a fair representation of other machine-learning-based severe weather models that would be used in an ensemble prediction setup. By comparing its performance with the two CNN-based methods, the contribution of CNN-based severe weather predictions can be examined.
@@ 3.4 Verification methods
To assess forecast quality from the three different ML-based prediction systems, we verify the ensemble mean severe weather probability against SPC reports from 1 January 2021–31 December 2021 using Brier Score (BS) and Brier Skill Score (BSS; Murphy (1973)). The climatology reference of BSS was derived from SPC reports between 1986 and 2015, and separately for locations, day of the year, and hour of the day. The spatial or temporally aggregated BSSs were computed as follows: given gridded severe weather probability values on the CONUS 80-km domain, the BSs were computed first for individual initialization days, forecast lead times, and grid cells. Then, the resulting three-dimensional arrays were averaged temporally or spatially. Finally, the climatology reference was averaged in the same way to produce BSSs. The above steps follow the suggestion of Hamill and Juras (2006). Three-component decomposition of BSs and reliability diagrams are also computed to attribute the BSS difference; their computation follows Murphy (1973) and Hsu and Murphy (1986).
@@ 4.1 Case-based assessments
Post-processing methods take HRRRv4 forecasts on 0200 and 0300 UTC 4 May 2021 as inputs and produce 4-hr, 80-km severe weather probabilities for $0$-0400 UTC ($0$ and 0100 UTC forecasts are ignored because of spin up. See Section 3). For CONUS grid cells that contain high UH HRRR forecasts, all three methods agreed well, with probabilities up to 60% (Fig. 5a–c).
@@ 4.2 Brier skill scores and reliability diagrams
The BSSs were aggregated over the CONUS for each 4-hr forecast lead time window (7). All three methods exhibited similar BSS variations (7a). BSSs were largest between 00-06Z, with the CGAN ensemble reaching 0.2 and the two baselines staying around 0.16. From 12-18Z, the BSSs of all methods decreased to their day-1 minimums, ranging from 0.04 to 0.06. Beyond 18Z, the BSSs of all three methods increased, with the CGAN ensemble mean approaching 0.15 and the two baselines near 0.12. In addition, the ensemble mean of all methods showed higher BSSs compared to most of the individual members, indicating that the ensemble ML predictions of severe weather was beneficial for improving forecast skills from deterministic predictions (7a).
All three methods generated forecasts with relatively low skill during the overnight through early morning hours (i.e., 09–15Z; (7a)). Several factors could be contributing to the relatively poor forecast skill. Fewer severe weather events are typically reported overnight, and thus, less training information is available. Further, overnight severe weather could be related to nocturnal convection that is often more challenging to forecast (e.g. Johnson and Wang, 2017; Blake et al., 2017). Errors in the input HRRR predictions may lead to poor post-processed forecasts, especially if errors are not biased in a systematic way that can be corrected by the ML algorithms.
For other forecast lead times, the CGAN ensembles outperformed the two baselines with statistically significant BSSs increases (7b,c) . The performance gains from the CNN baseline to the CGAN ensemble indicate that the CGANs contributed positively to the CNN-based prediction model. The CNN baseline and MLP baseline showed comparable BSS performance, with the CNN baseline being slightly better, especially for 06-12Z (7a).
Figure 8: (a) Grid-point-wise BSSs computed using the CGAN ensemble mean for all initialization days and forecast lead times. (b) As in (a), but for the CNN ensemble mean. (c) As in (a) but for the MLP ensemble. (d) is the difference between (a) and (b). (e) is the difference between (a) and (c). Neighboring grid cells are included in the computation of grid-point-wise BSSs. Grid cells with severe weather reports lower than 150 are marked using hatches.
@@
Note that the CGAN ensemble performed worse than the two baselines in some grid cells. This was primarily due to the positional error of the HRRR forecasts. One of these cases is shown in Section 4.4.1. We think that with CAM forecast skill being improved in the future, such double penalty problems will be substantially reduced.

### v2

@@ Abstract
[…]
We use a new machine learning (ML) technique to generate probabilistic forecasts of convective weather hazards, such as tornadoes and hail storms, with the output from high-resolution numerical weather model forecasts. The new ML system generates an ensemble of synthetic forecast fields from a single forecast, which are then used to train ML models for convective hazard prediction. Using this ML-generated ensemble for training leads to improvements of 10–20% in severe weather forecast skills compared to using other ML algorithms that use only output from the single forecast. This work is unique in that it explores the use of ML methods for producing synthetic forecasts of convective storm events and using these to train ML systems for high-impact convective weather prediction.
@@ 1 Introduction
Validating the probabilistic severe weather predictions and exploring the advances of deep generative models are the main motivations of this work. Ensemble predictions have clear benefits. The best guess of an ensemble is typically more skillful than deterministic predictions. The ensemble spread measures the predictive uncertainties; it answers the question of “how much can we trust the predicted results?”. Given the difficulty and social impact of severe weather post-processing, organizing it within an ensemble prediction workflow is essential. Many deterministic, point-based machine learning methods can make ensemble predictions if they have an ensemble of CAM runs as inputs. However, maintaining and pre-processing a large ensemble of CAMs may introduce technical challenges such as data transmission and storage. The development of deep generative models provides an alternative. In this research, deep generative models are trained to create an ensemble of synthetic forecasts from a deterministic CAM run. This synthetic ensemble will provide convective-scale information and support the estimation of severe weather. Deep generative models are neural networks that can approximate probability distributions, and notably, they can create synthetic outputs from the data distributions they have learned (Creswell et al., 2018; Goodfellow et al., 2014; Ruthotto and Haber, 2021). Deep generative models have achieved success in downscaling [e.g., Leinonen et al. (2020) for satellite imagery, Miralles et al. (2022) for surface wind], nowcasting [e.g., Ravuri et al. (2021), Luo et al. (2022), and Tian et al. (2019) for precipitation nowcasting, Gong et al. (2023) for radar reflectivity], and data-driven ensemble weather forecasting [Zhong et al. (2023) for FuXi-Extreme, Price et al. (2023) for GenCast] problems, indicating that they can be applied to generate high-resolution meteorological fields.
In this research, a novel ML-based CAM post-processing method is introduced that combines a CNN-based prediction model and a deep generative model that creates synthetic CAM forecasts as input predictors for severe weather post-processing. The generative model can effectively generate an ensemble of severe weather predictions from a deterministic CAM run, potentially improving forecast skills and uncertainty estimates. The post-processing system uses CAM output on the native model grid, eliminating the need for upscaling that has been used in prior work (e.g. Loken et al., 2020; Sobash et al., 2020). The system is trained to output probabilities of severe weather at grid points within the CAM domain using observations of severe weather events.
@@ 2.2 Forecast data
We used both HRRR version 3 and 4 (Dowell et al., 2022). The HRRR is an operational, real-time CAM based on WRF-ARW [Weather Research and Forecasting model, WRF; Advanced Research WRF, ARW; Skamarock et al. (2019)]. HRRR has 3 km horizontal grid spacing and 50 vertical levels; it takes Rapid Refresh (RAP; Benjamin et al. (2016)) as initial and boundary conditions, producing hourly, deterministic forecasts over CONUS. Here, the $0$ UTC initializations were used, with HRRRv3 covering 12 July 2018 to 2 December 2020, and HRRRv4 afterwards. In addition, experimental HRRRv4 initialized from 1 October 2019 to 2 December 2020 was also used.
@@ 2.3 Observations
Severe weather reports collected by the SPC at the National Oceanic and Atmospheric Administration (hereafter, “SPC reports”) were used as the observational target. SPC reports provide the location and time for tornadoes, hail, and convective wind gusts that have sufficient intensities and societal impact. Such information is placed on the gridded 80-km CONUS domain by matching the closest grid cell to the starting location of each report. For temporal dimensions, SPC reports were aggregated hourly with a 4-hour time window. That said, given a severe weather report on hour $t$; it will be mapped to the 80-km CONUS domain four times, on $t-2$, $t-1$, $t$, and $t+1$. This choice is based on the operational demand for 4-hour severe weather outlooks (e.g. Krocak and Brooks, 2020).
Gridded SPC reports were used for the training and verification of the post-processing models, as well as the estimation of climatology references (Fig. 1.b and d). We prefer the gridded SPC reports over the original ones because the spatiotemporal re-gridding creates a “tolerance” for the regional and population bias that SPC reports may potentially have (e.g. Doswell et al., 1999). Hereafter, when “SPC reports” is mentioned, it means the re-gridded version on the 80-km CONUS domain.
@@ 3 Methods
The above post-processing method was trained from 5 July 2018 to 31 December 2020 with HRRRv3 and HRRRv4 forecasts as inputs and SPC reports as targets. It produces 4-hr, 80-km severe weather probabilities out to 24 hours with the exceptions of 00Z and 01Z; these two forecast lead times were ignored because of the HRRR model spin-up. The validation set was a 10% random split from the training set, and it was fixed for all the training steps. The verification period of the final post-processing outputs is 1 January 2021–31 December 2021 and with HRRRv4 only.
@@ 3.1 Conditional Generative Adversarial Networks
Figure 3: (a) The training procedure of CGAN. (b) The architecture of the CNN-based generator; it contains convolutional layers (“conv”), Rectified Linear Unit (ReLU) activation function, and Batch Normalization (BN). Numbers beside each block represent the number of convolution kernels. (c) similar to (b), but for the architecture of the discriminator. Note that the CNN-based generator (b) will be preserved after training, whereas the discriminator is used for the CGAN training only.
This research applies CGANs to generate synthetic ensembles from deterministic HRRR forecasts. CGANs are Generative Adversarial Networks (GANs) that utilize conditional information as additional inputs to enable more control over the sample generation process (Mirza and Osindero, 2014; Isola et al., 2017). Here the conditional information is provided by five storm-scale explicit predictors: 2-5 km UH, 0-2 km UH, APCP, GRPL, and 10-m SPD. By accepting such conditional inputs, CGANs were trained to generate synthetic CREF and environmental predictors. The idea behind this choice is that UH and CREF are the two major CAM outputs that are related to convective storms directly (Clark et al., 2012; Gallo et al., 2021). If both UH and CREF are allowed to be generated by CGANs freely, the resulting synthetic ensembles may contain unrealistic storms and bias the severe weather prediction step. Thus, training CGANs to create CREF, but letting the generation process be conditioned on UH and other explicit predictors would balance the creativity and realism of CGAN outputs.
The CGAN in this research consists of a CNN-based generator and discriminator (Fig. 3a). The generator has two input branches, one takes conditional inputs, and the other accepts the initial state of predictors to be generated. Such initial states were prepared by combining the original HRRR predictors with $\mathcal{N}(0,0.5)$ distributed Gaussian noise (Fig. 3b). Negative values will be corrected to zeros for the initial states of CAPE, CIN, and CREF. The base architecture of the generator is an encoder-decoder CNN with skip connections on each encoding level, known as UNET (Ronneberger et al., 2015). The UNET applies convolutional layers with two strides and transposed convolutional layers for down- and up-samplings, respectively. The number of convolution kernels doubles after each down-sampling operation, from 32 to 512, and the numbers are symmetrical on the upsampling side (Fig. 3b). The discriminator of the CGAN is a CNN-based binary classifier that aims to distinguish generator outputs from their corresponding HRRR predictors. Besides the classification inputs, the discriminator also takes the same conditional inputs as that of the generator (Fig. 3c).
@@
Two CGANs of the above were proposed: one generates CAPE, CIN, and CREF (i.e., three input channels), and the other generates MSLP, 2-m Temp, 2-m Dewpoint T, 0-1 km SRH, 0-3 km SRH, 0-6 km U Shear, 0-6 km V Shear (i.e., seven input channels). The same CGANs were applied to all forecast lead times.
@@ 3.2 CNN-based end-to-end severe weather prediction
Figure 4: (a) The CNN-based end-to-end severe weather prediction model. (b) the architecture of the 2-d CNN in (a); it contains convolutional layers (“conv”), Rectified Linear Unit (ReLU) activations, and Batch Normalization (BN). (c) The architecture of the classification model in (a). (d) The architecture of the MLP-based severe weather prediction model. Monte Carlo (MC) dropout layers are included in (c) and (d)
@@
The representation learning model was trained from 5 July 2018 to 31 December 2020, using HRRRv3 and v4 forecasts. SPC reports on the corresponding forecast lead times were used as training targets. This is achieved by configuring an additional dense layer with a sigmoid activation function to compute cross-entropy loss (Fig. 4b).
3.3 Hyperparameter optimization and baseline methods
@@
The MLP ensemble does not rely on CNNs to make severe weather predictions. Instead, it combines an MLP with feature engineering steps. Given the same 64-by-64-sized HRRR predictors that CNN-based methods would use, the MLP baseline converts them into scalar inputs by computing their mean and maximum values on spatial and forecast lead time dimensions. This pre-processing step is similar to Sobash et al. (2020). The architecture of MLP is illustrated in Fig. 4d; it consists of two hidden layers with BN, Rectified Linear Unit (ReLU) activation function, and MC dropout. Several existing works have shown that MLPs are effective in deriving severe weather probabilities from CAM outputs (e.g. Sobash et al., 2020). Notably, they can outperform the conventionally used surrogate methods. Thus, in this experiment, the MLP baseline is a fair representation of other machine-learning-based severe weather models that would be used in an ensemble prediction setup. By comparing its performance with the two CNN-based methods, the contribution of CNN-based severe weather predictions can be examined.
Hyperparameter optimizations were conducted on the main method and the two baselines. For each neural network, the relative importance of its hyperparameter options was identified based on the prior knowledge of the network design and other existing studies (e.g. Greff et al., 2016; Sharma et al., 2019; Yu and Zhu, 2020). Experiment trials were then proposed from the most important to the least important options. The performance criteria of the CGAN trials were the reconstruction loss and stability. Low reconstruction loss across all the generated predictors was preferred. For the CNN-based representation learning model, CNN-based classifier, and the MLP baseline, their performance criteria were severe weather prediction skills averaged across all forecast lead times. The hyperparameter optimization of this study was conducted based on the same training and validation set, and without accessing the verification data. For the two CGANs, each trial was fixed to 200 epochs. For the CNN-based representation learning model, CNN-based classifier, and the MLP baseline, their trials had early stopping implemented. At the end of the hyperparameter search, neural network weights from the best trials were preserved.
The results of hyperparameter optimization for all models, from the most important to the least important options are provided as follows:
  • CGAN: $\mathcal{N}(0,0.5)$ additive noise, 1e-4 learning rate with decay, batch size 32, Adam optimizer (Kingma and Ba, 2014), ReLU activation (see Fig. 3b).
  • CNN-based representation learning model: sample rebalancing of 1 positive: 10 negatives, 1e-4 learning rate with decay, batch size 32, Adam optimizer, ReLU activation, {48, 64, 96, 128} number of convolution kernels and 2 convolutional layer stacks per downsampling block, global max pooling (see Fig. 4b).
  • CNN-based classifier: sample rebalancings of 1:1, dropout rate 0.1, 1e-4 learning rate with decay, batch size 32, Adam optimizer, ReLU activation, 128 1-d convolution kernels and 64 dense layer neurons (see Fig. 4c).
  • MLP baseline: sample rebalancings of 1:1, dropout rate 0.1, 1e-4 learning rate with decay, batch size 64, Adam optimizer, ReLU activation, 128, and 64 neurons for the two dense layers (see Fig. 4d).
The model training and hyperparameter optimizations of this study were conducted on NVIDIA Tesla V100 GPUs with additional CPUs to support the data pipeline. The MLP baselines were roughly 30% faster than the CNN baseline and 70% faster than the CGAN method.
@@ 3.4 Verification methods
To assess forecast quality from the three different ML-based prediction systems, we verify the ensemble mean severe weather probability against SPC reports from 1 January 2021–31 December 2021 using Brier Score (BS) and Brier Skill Score (BSS; Murphy (1973)). The climatology reference of BSS was derived from SPC reports between 1986 and 2015, and separately for locations, day of the year, and hour of the day. The spatial or temporally aggregated BSSs were computed as follows: given gridded severe weather probability values on the CONUS 80-km domain, the BSs were computed first for individual initialization days, forecast lead times, and grid cells. Then, the resulting three-dimensional arrays were averaged temporally or spatially. Finally, the climatology reference was averaged in the same way to produce BSSs. The above steps follow the suggestion of Hamill and Juras (2006). Three-component decomposition of BSs and reliability diagrams were also computed to attribute the BSS difference; their computation follows Murphy (1973) and Hsu and Murphy (1986).
@@ 4.1 Case-based assessments
Post-processing methods take HRRRv4 forecasts on 0200 and 0300 UTC 4 May 2021 as inputs and produce 4-hr, 80-km severe weather probabilities for $0$-0400 UTC ($0$ and 0100 UTC forecasts were ignored because of spin up. See Section 3). For CONUS grid cells that contain high UH HRRR forecasts, all three methods agreed well, with probabilities up to 60% (Fig. 5a–c).
@@ 4.2 Brier skill scores and reliability diagrams
The BSSs were aggregated over the CONUS for each 4-hr forecast lead time window (Fig. 7). All three methods exhibited similar BSS variations (Fig. 7a). BSSs were largest between 00-06Z, with the CGAN ensemble reaching 0.2 and the two baselines staying around 0.16. From 12-18Z, the BSSs of all methods decreased to their day-1 minimums, ranging from 0.04 to 0.06. Beyond 18Z, the BSSs of all three methods increased, with the CGAN ensemble mean approaching 0.15 and the two baselines near 0.12. In addition, the ensemble mean of all methods showed higher BSSs compared to most of the individual members, indicating that the ensemble ML predictions of severe weather was beneficial for improving forecast skills from deterministic predictions (Fig. 7a).
All three methods generated forecasts with relatively low skill during the overnight through early morning hours (i.e., 09–15Z; (Fig. 7a)). Several factors could be contributing to the relatively poor forecast skill. Fewer severe weather events are typically reported overnight, and thus, less training information is available. Further, overnight severe weather could be related to nocturnal convection that is often more challenging to forecast (e.g. Johnson and Wang, 2017; Blake et al., 2017). Errors in the input HRRR predictions may lead to poor post-processed forecasts, especially if errors are not biased in a systematic way that can be corrected by the ML algorithms.
For other forecast lead times, the CGAN ensembles outperformed the two baselines with statistically significant BSSs increases (Fig. 7b,c) . The performance gains from the CNN baseline to the CGAN ensemble indicate that the CGANs contributed positively to the CNN-based prediction model. The CNN baseline and MLP baseline showed comparable BSS performance, with the CNN baseline being slightly better, especially for 06-12Z (Fig. 7a).
Figure 8: (a) Grid-point-wise BSSs computed using the CGAN ensemble mean for all initialization days and forecast lead times. (b) As in (a), but for the CNN ensemble mean. (c) As in (a) but for the MLP ensemble mean. (d) is the difference between (a) and (b). (e) is the difference between (a) and (c). Neighboring grid cells are included in the computation of grid-point-wise BSSs. Grid cells with severe weather reports lower than 150 are marked using hatches. (f) The categorical distributions of the top 11.45% most successfully predicted severe weather cases from the CGAN ensemble mean, the CNN ensemble mean, and the MLP ensemble mean.
@@
Note that the CGAN ensemble performed worse than the two baselines in some grid cells. This was primarily due to the positional error of the HRRR forecasts. One of these cases is shown in Section 4.4.1. We think that with CAM forecast skills being improved in the future, such double penalty problems will be substantially reduced.
The performance of severe weather prediction models varies by both regions and severe weather categories. In Fig. 8f, the categorical distributions of correctly predicted severe weather events were provided for each method. The purpose of this analysis is to investigate which severe weather categories are more challenging to predict for each post-processing method. For severe weather events within the verification data, their individual Brier scores (i.e., the squared difference between the predicted probability and 1.0) associated with the ensemble mean of each post-processing method were ranked, with lower scores indicating more successful predictions. In this analysis, the top 10% of the most successful predictions of each method were selected for each forecast lead time, and their severe weather categories were summarized.
From Fig. 8, all methods tended to perform well in combined severe weather events. This is likely because these combined events have left stronger signals within the HRRR forecasts. For example, given 711 combined tornado and wind gust events, all methods captured at least 210 of them with good Brier scores. The combined tornadoes and wind gusts were likely caused by supercell thunderstorms; they created an ideal environment for tornado genesis and brought wind gust damage from downbursts. Supercell thunderstorms have clear connections with UH, and thus, the resulting severe weather events were predicted well by all methods. On the contrary, isolated events, especially “tornadoes only” and “hail only,” were less likely to achieve the top 10% Brier scores. These isolated events could be relatively mild (i.e., non-supercell tornadoes) or occurred in regions where HRRR forecast skills were poor (e.g., hail reports on the West Coast and the Rockies), thus leading to somewhat unsatisfactory predictions. In summary, the predictability of each severe weather category has a larger impact than the choice of post-processing methods. Combined severe weather events (e.g., tornadoes and wind gusts) are more likely to be predicted successfully.